two train 120m and 80m in length are running in opposite direction with velocity 40 km/hr and 30km/hr. in what time they will completely cross each other
Answer by mananth(16946)

two train 120m and 80m in length are running in opposite direction with velocity 40 km/hr and 30km/hr. in what time they will completely cross each other

the sum of lengths of the trains is the distance
120 + 80 = 200 m
=0.2 km
the trains are in opposite direction so effective speed = 40+30=70 km/h
t= d/r
t=0.2/70
t=0.00286 hours
1 hour = 3600 seconds
0.00286 * 3600 = 10.29 seconds